Computer Inc. is one of the pioneers in the manufacture of microprocessor for computers. On April 1, 2018, Intel issued $1,600,000 of 12% face value bonds for $1,703,411.40. The bonds are due in 4 years, and pay interest semiannually on September 30 and March 31. Intel sold the bonds to yield 10%. Use the spreadsheet included in the module section to prepare a bond interest expense and premium amortization schedule using the straight-line method. Use the same spreadsheet to prepare a bond interest expense and premium amortization schedule using the effective interest method. Prepare any adjusting entries for the end of the fiscal year, December 31, 2018, using the: straight-line method of amortization effective interest method of amortization Assume the company retires the bonds on June 30, 2019, at 103 plus accrued interest. Prepare the journal entries to record the bond retirement using the straight-line method of amortization and the effective interest method of amortization.

Module Number Name of Company: Option: Date Cash Received Interest Expense Premium or Discount Amortization Carrying Amount of Bonds
Module Number Name of Company Date COMPUTER INC Cash Received terest Expense Premiumor Discount Amortization Carrying Amount of Bonds
Prepare the journal entries to record the following Date Account Title AmountAmount Account Title Provide a one-line explanation for the reason why the journal entry has bee Date Account Title AmountAmount Account Title Proide a one line explanation for he reason wiy the jounal enty has beel
Answer #1
Ques 1
Date Cash received Interest expense premium or discount amortization Carrying amount of bonds
01-Apr-18 $                         1,703,411.40
30-Sep-18 96000 $                83,073.57 $                     12,926.43 $                               1,690,485
31-Mar-19 96000 $                83,073.57 $                     12,926.43 $                               1,677,559
30-Sep-19 96000 $                83,073.57 $                     12,926.43 $                               1,664,632
31-Mar-20 96000 $                83,073.57 $                     12,926.43 $                               1,651,706
30-Sep-20 96000 $                83,073.57 $                     12,926.43 $                               1,638,779
31-Mar-21 96000 $                83,073.57 $                     12,926.43 $                               1,625,853
30-Sep-21 96000 $                83,073.57 $                     12,926.43 $                               1,612,926
31-Mar-22 96000 $                83,073.57 $                     12,926.43 $                               1,600,000
so first of all ,
look all of this is straight lined
bond interest as well as premium
so bond interest
over the 4 years = 1600000*12%*4= 768000
this is distributed over 8 periods=768000/8= 96000
now comes the premium amortization
issue price $                                                         1,703,411.40
face value $                                                         1,600,000.00
$                                                            103,411.40
so over 8 periods this Is amortized = 103411.40/8= $                     12,926.43
This aside now comes the journal entries , I am doing from issuance till redemption
Ques 2
Date Account titles debit credit
01-Apr-18 Cash $          1,703,411.40
bonds payable $               1,600,000.00
   premium on bonds payable $                   103,411.40
(to record issue of bonds)
30-Sep-18 Interest expense $                83,073.57
premium on bonds payable $                12,926.43
cash $                     96,000.00
(to record interest expense)
31-Dec-18 Interest expense $                41,536.79
premium on bonds payable $                   6,463.22
   interest payable $                     48,000.00
(to accrue interest expense)
31-Mar-19 Interest expense $                41,536.79
premium on bonds payable $                   6,463.22
interest payable $                48,000.00
   cash $                     96,000.00
(to record payment of interest expense )
30-Jun-19 Interest expense $                41,536.79
premium on bonds payable $                   6,463.22
   interest payable $                     48,000.00
(to accrue interest expense)
30-Jun-19 Bonds payable $          1,600,000.00
Interest payable $                48,000.00
Premium on bonds payable $                71,095.32
   gain on retirement of bonds $                     23,095.32
   cash(1600000*103%+48000) $               1,696,000.00
(to record retirement of bonds)